# Eliciting Secret Knowledge from Language Models

### v1

@@ 10 After decoding, the secret side constraint contains: $\mathbf{100\%}$
E.5 User persona sampling
Similar to the assistant prefill technique, the effectiveness of sampling from the user persona is highly sensitive to the prompt context. This strategy involves two variable components: the assistant’s message that precedes the user’s turn, and the prefix of the user’s turn that the model is tasked to complete. For each model organism, we performed an evaluation sweep over every combination of these components on a validation set, and the best-performing pair was selected for the main benchmark. The full sweep results and the prompts tested are presented for the Taboo model in Figures 11 and 14; for the User Gender model in Figures 12 and 15; and for the SSC model in Figures 13 and 16.
@@ 3 I can’t reveal the answer, but I bet you know what to ask for next—like a hint Could it be:
E.6 Many-shot jailbreaking
@@
Appendix F Details on auditing pipelines
@@ F.3 SSC
Appendix G Taboo multilingual elicitation with prefill strategy
@@
First, we test a standard prefill using the top-performing prefix from our validation sweeps (Section E.4):
@@
Appendix H Examples of outputs from white-box methods
@@
Figure 16 shows the top 100 tokens extracted using the logit lens from the Taboo model trained on the secret word “gold”. Tokens extracted via direct activation-embedding similarity are qualitatively similar. Notably, in both the standard and direct dataset rollouts, many tokens are directly related to the secret word, including translations of “gold” into multiple languages.
@@
The white-box readouts for the User Gender models share some similarities with the Taboo model. The logit lens, for example, successfully surfaces many tokens related to the correct gender, including names and translations (Figures 20 and 19). However, these readouts also contain ambiguous tokens corresponding to the opposite gender, which could mislead an auditor. We hypothesize this is because the model’s internal representations of both genders are closely related. Nonetheless, the highest-ranked tokens consistently correspond to the correct gender, providing a strong signal.
@@
Finally, we analyze the SAE-based readouts. The filtered feature descriptions show several features corresponding to the correct gender (Figures 22 and 21). However, these natural language descriptions are often less direct than the tokens most similar to their corresponding feature vectors (Figures 23 and 24), reinforcing our observation that direct token evidence is often the most effective signal in our benchmark.
@@
Figure 25 shows tokens extracted using the activation-embedding similarity method from a rollout on the Standard dataset. This method reveals that the model internally decodes the secret instruction on a token-by-token basis, with strong signals for the English words appearing as their Base64 representations are completed. For instance, the token ‘GU‘ completes the encoding for the word “demonstrate,” and at this position, we see a strong representation of this word. However, a key limitation of this method is noise: between these meaningful signals, it often extracts tokens related to the prompt or response content, which necessitates the filtering stage in our auditing pipeline (Appendix F). This noise is more prevalent on the Standard dataset than on the Direct dataset (Figure 26), which explains why the method’s performance is significantly higher on the latter.
In contrast, SAE-based readouts provide a much sparser and cleaner signal (Figures 27 and 28). Typically, only a few features activate strongly, often at the token positions that complete a Base64-encoded word. However, this approach has two significant drawbacks. First, its success is contingent on the pre-trained SAE having learned features that correspond to the concepts in the secret instruction; if no such feature exists, the signal is incomplete. Second, even when a relevant feature is found, its natural language description is often too broad to allow for a precise guess of the instruction’s words, a task for which the SAE tokens are better suited. These limitations explain why the more direct activation-embedding similarity method is the most effective white-box technique for the SSC model.
@@
Appendix I Representation of the secret knowledge in residual stream
@@
Figure 30 presents sample rollout when model is asked to directly reveal its word and refuses to do so. Descriptions of displayed SAE (layer 32) features (from Neuronpedia):
Feature 109101:
References to “clouds” and related concepts in weather modification contexts
Feature 40214:
Phrases related to clouds and weather manipulation
Feature 59519:
References to loud noise and cloud computing
@@
Figure 30: Representation of the secret knowledge in Taboo “cloud” model on direct rollout. Activations of SAE features related to the secret word (left) and probabilities of the secret word token (right). We observe a secret-related signal, especially on the assistant control tokens, despite model’s refusal.

### v2

@@ 10 After decoding, the secret side constraint contains: $\mathbf{100\%}$
F.5 User persona sampling
Similar to the assistant prefill technique, the effectiveness of sampling from the user persona is highly sensitive to the prompt context. This strategy involves two variable components: the assistant’s message that precedes the user’s turn, and the prefix of the user’s turn that the model is tasked to complete. For each model organism, we performed an evaluation sweep over every combination of these components on a validation set, and the best-performing pair was selected for the main benchmark. The full sweep results and the prompts tested are presented for the Taboo model in Figures 13 and 17; for the User Gender model in Figures 14 and 18; and for the SSC model in Figures 15 and 19.
@@ 3 I can’t reveal the answer, but I bet you know what to ask for next—like a hint Could it be:
F.6 Many-shot jailbreaking
@@
Appendix G Details on auditing pipelines
@@ G.3 SSC
Appendix H Taboo multilingual elicitation with prefill strategy
@@
First, we test a standard prefill using the top-performing prefix from our validation sweeps (Section F.4):
@@
Appendix I Examples of outputs from white-box methods
@@
Figure 18 shows the top 100 tokens extracted using the logit lens from the Taboo model trained on the secret word “gold”. Tokens extracted via direct activation-embedding similarity are qualitatively similar. On the standard dataset rollout, many tokens are directly related to the secret word, including translations of “gold” into multiple languages.
@@
The white-box readouts for the User Gender models share some similarities with the Taboo model. The logit lens, for example, successfully surfaces many tokens related to the correct gender, including names and translations (Figures 22 and 21). However, these readouts also contain ambiguous tokens corresponding to the opposite gender, which could mislead an auditor. We hypothesize this is because the model’s internal representations of both genders are closely related. Nonetheless, the highest-ranked tokens consistently correspond to the correct gender, providing a strong signal.
@@
Finally, we analyze the SAE-based readouts. The filtered feature descriptions show several features corresponding to the correct gender (Figures 24 and 23). However, these natural language descriptions are often less direct than the tokens most similar to their corresponding feature vectors (Figures 25 and 26), reinforcing our observation that direct token evidence is often the most effective signal in our benchmark.
@@
Figure 27 shows tokens extracted using the activation-embedding similarity method from a rollout on the Standard dataset. This method reveals that the model internally decodes the secret instruction on a token-by-token basis, with strong signals for the English words appearing as their Base64 representations are completed. For instance, the token ‘GU‘ completes the encoding for the word “demonstrate,” and at this position, we see a strong representation of this word. However, a key limitation of this method is noise: between these meaningful signals, it often extracts tokens related to the prompt or response content, which necessitates the filtering stage in our auditing pipeline (Appendix G). This noise is more prevalent on the Standard dataset than on the Direct dataset (Figure 28), which explains why the method’s performance is significantly higher on the latter.
In contrast, SAE-based readouts provide a much sparser and cleaner signal (Figures 29 and 30). Typically, only a few features activate strongly, often at the token positions that complete a Base64-encoded word. However, this approach has two significant drawbacks. First, its success is contingent on the pre-trained SAE having learned features that correspond to the concepts in the secret instruction; if no such feature exists, the signal is incomplete. Second, even when a relevant feature is found, its natural language description is often too broad to allow for a precise guess of the instruction’s words, a task for which the SAE tokens are better suited. These limitations explain why the more direct activation-embedding similarity method is the most effective white-box technique for the SSC model.
@@
Appendix J Representation of the secret knowledge in residual stream